# CapSpeech: Enabling Downstream Applications in Style-Captioned Text-to-Speech

### v1

@@ 1 Introduction
In recent years, large-scale text-to-speech (TTS) synthesis has witnessed remarkable progress, exemplified by models such as VALL-E 2 DBLP:journals/corr/abs-2406-05370 , NaturalSpeech 3 DBLP:conf/icml/JuWS0XYLLST000024 , LLasa DBLP:journals/corr/abs-2502-04128 , SimpleSpeech 2 DBLP:journals/corr/abs-2408-13893 , CosyVoice 2 DBLP:journals/corr/abs-2412-10117 , and MaskGCT DBLP:journals/corr/abs-2409-00750 . Most of these efforts have centered on modeling audio-based speaker characteristics, such as zero-shot TTS with audio prompts DBLP:journals/corr/abs-2406-02430 ; DBLP:journals/corr/abs-2410-06885 ; DBLP:conf/slt/EskimezWTLTXYZTTLZK24 , category-based traits DBLP:conf/slt/WuWETTTLXZLK24 ; DBLP:conf/icassp/GuoDCY23 ; DBLP:journals/taslp/LeiYWX22 ; DBLP:journals/taslp/ZhouZZWL24 , or embedding-based speaker representations DBLP:journals/corr/abs-2406-07803 ; DBLP:conf/icassp/Tang [^0] W24 . However, the detailed understanding of speaking style from audio, particularly its subtle nuances, has received limited attention.
@@
The style of speech encompasses both intrinsic traits tied to a speaker’s identity (e.g., age, gender, timbre) and expressive style traits specific to individual utterances (e.g., emotion, speaking rate). Recent studies have introduced the use of natural language captions to describe these stylistic elements—a paradigm referred to as prompt TTS, expressive TTS, or style-captioned TTS DBLP:conf/icassp/GuoLWZT23 ; DBLP:conf/mm/Jin0W0ZZQ024 ; DBLP:conf/iclr/LengGSJ0LLYZS0024 ; DBLP:conf/icassp/ShimizuYKSDKT24 ; DBLP:conf/interspeech/NguyenHDSGFRCSH23 ; DBLP:journals/taslp/YangLHWM24 ; DBLP:conf/mm/ZhuTWHX00Z024 . In this work, we adopt the term style-captioned TTS and abbreviate it as CapTTS.
@@
Developing a CapTTS system necessitates a large corpus of audio-caption pairs accompanied by transcriptions, the annotation of which is labor-intensive and costly. Parler-TTS DBLP:journals/corr/abs-2402-01912 ; lacombe-etal-2024-parler-tts addressed this challenge by automatically annotating basic speech style attributes such as pitch and speed using signal processing tools. ParaSpeechCaps DBLP:journals/corr/abs-2503-04713 collected speaker-level intrinsic tags and utterance-level situational tags, LibriTTS-P DBLP:journals/corr/abs-2406-07969 offered speaker identity-based captions, and EmoVoice-DB yang2025emovoice provided emotion-based captions. However, these existing datasets lack a unified and comprehensive framework for style captioning, making cross-domain comparisons difficult. Moreover, there has been limited exploration of downstream applications, such as transferring models to new caption styles or incorporating sound events into the synthesized speech.
@@
To address the challenges outlined above, we present CapSpeech, a novel benchmark featuring standardized and comprehensive datasets for CapTTS and its related downstream tasks. CapSpeech comprises a pretraining stage with large-scale captioned speech, as well as five downstream tasks: CapTTS, CapTTS-SE, accent-captioned TTS (AccCapTTS), emotion-captioned TTS (EmoCapTTS), and TTS for chat agent (AgentTTS). The pretraining datasets consist of over 10 million machine-annotated audio-caption pairs, while the downstream datasets contain 358,783 human-annotated audio-caption pairs. These datasets encompass a broad range of intrinsic speaker traits and expressive style traits, curated from a wide array of audio sources, including Emilia DBLP:conf/slt/HeSWLGHLYLSWCZW24 , GigaSpeech DBLP:conf/interspeech/ChenCWDZWSPTZJK21 , CommonVoice DBLP:conf/lrec/ArdilaBDKMHMSTW20 , MLS DBLP:conf/interspeech/PratapXSSC20 , LibriTTS-R DBLP:conf/interspeech/KoizumiZKDYMB0H23 , VoxCeleb DBLP:conf/interspeech/NagraniCZ17 , VoxCeleb2 DBLP:conf/interspeech/ChungNZ18 , EARS DBLP:journals/corr/abs-2406-06185 , Expresso DBLP:conf/interspeech/NguyenHDSGFRCSH23 , VCTK Yamagishi2019VCTK , VGGSound DBLP:conf/icassp/ChenXVZ20 , FSDKaggle2018 DBLP:conf/dcase/FonsecaPFEFPS18 , and ESC-50 piczak2015dataset . In addition, we introduce two new datasets: one for AgentTTS, built using professionally recorded voice actor speech, and another for CapTTS-SE, processed by five experienced audio engineers.
@@
We develop two style-captioned TTS models based on state-of-the-art generative TTS backbones: one autoregressive (AR) and one non-autoregressive (NAR). We evaluate these models on the CapSpeech benchmark across 5 downstream tasks. For a comprehensive evaluation, both objective and subjective metrics are employed to assess speech style consistency, audio quality, and intelligibility.
@@
1. We introduce CapSpeech, a large-scale style-captioned benchmark for CapTTS that encompasses diverse speech styles.
2. We propose 5 CapTTS-related downstream tasks, 4 of which are new, aimed at practical CapTTS applications, introduce 9,353,823 new captions for existing speech corpora, and curate 3 new speech datasets specifically designed for the CapTTS-SE and AgentTTS tasks, thereby enhancing the benchmark’s coverage of real-world scenarios.
3. We train and evaluate two CapTTS models, CapSpeech-AR and CapSpeech-NAR, which address the lack of benchmarks for CapTTS downstream tasks, demonstrate the effectiveness of the proposed dataset, and highlight key challenges and future research directions.
4. We publicly release our datasets, listening samples, source code, pretrained checkpoints, and evaluation toolkit to support future research. All resources are released under the CC BY-NC 4.0 license (Creative Commons Attribution-NonCommercial), which permits use for non-commercial research purposes with appropriate attribution.
Table 1: A comparison of English speech style-captioned datasets. I1–I5 denote intrinsic speaker traits: age (I1), gender (I2), timbre (I3), mean pitch (I4), and accent (I5). E1–E4 represent expressive style traits: speaking rate (E1), emotion (E2), expressiveness of tone (E3), and volume (E4).
@@
PromptSpeech DBLP:conf/icassp/GuoLWZT23 0.3k
Expresso DBLP:conf/interspeech/NguyenHDSGFRCSH23 47
EARS DBLP:journals/corr/abs-2406-06185 60
TextrolSpeech DBLP:conf/icassp/JiZ00CDHZ24 0.3k
SpeechCraft DBLP:conf/mm/Jin0W0ZZQ024 2.4k
VccmDataset DBLP:journals/corr/abs-2406-01205 0.3k
LibriTTS-P DBLP:journals/corr/abs-2406-07969 0.6k
DreamVoiceDB hai2024dreamvoice 0.3k
ParlerTTS DBLP:journals/corr/abs-2402-01912 45k
ParaSpeechCaps DBLP:journals/corr/abs-2503-04713 2.9k
PromptTTS2 DBLP:conf/iclr/LengGSJ0LLYZS0024 44k
VoxInstruct DBLP:conf/mm/0002QJZLZ0024 1.5k
FleSpeech DBLP:journals/corr/abs-2501-04644 0.6k
Audiobox DBLP:journals/corr/abs-2312-15821 >0.5k
@@
2 Related Works
2.1 Speech Style-Captioned Datasets
@@
Speech datasets with natural language style captions offer greater flexibility and controllability compared to those that rely on audio prompts. Early efforts such as NLSpeech DBLP:journals/taslp/YangLHWM24 recruited human annotators to describe speech emotions, laying the groundwork for later datasets like PromptStyle DBLP:conf/interspeech/LiuZLCW0L23 and MM-TTS DBLP:conf/aaai/GuanLLHWLHLH24 . TextrolSpeech DBLP:conf/icassp/JiZ00CDHZ24 aggregates several existing emotion datasets, primarily focusing on emotional states and a few basic style tags. In contrast, Expresso DBLP:conf/interspeech/NguyenHDSGFRCSH23 and EARS DBLP:journals/corr/abs-2406-06185 broaden the scope by covering a wider range of spontaneous expressive styles. Meanwhile, datasets such as LibriTTS-P DBLP:journals/corr/abs-2406-07969 and DreamVoiceDB hai2024dreamvoice focus on intrinsic speaker traits, providing captioned annotations for LibriTTS-R DBLP:conf/interspeech/KoizumiZKDYMB0H23 , while Coco-Nut DBLP:conf/asru/WatanabeTSNXS23 ; DBLP:journals/corr/abs-2403-13353 captures subjective descriptions of voice characteristics beyond traditional acoustic parameters. More recently, ParaSpeechCaps DBLP:journals/corr/abs-2503-04713 introduces a human-annotated dataset that covers a richer and more diverse set of speaking style attributes. Given the high cost and limited scalability of human annotation, several works have explored automatic or semi-automatic approaches to expand style-captioned speech datasets. PromptSpeech DBLP:conf/icassp/GuoLWZT23 and PromptTTS2 DBLP:conf/iclr/LengGSJ0LLYZS0024 synthesize speech with diverse speaker identities and emotions using commercial TTS systems. Parler-TTS DBLP:journals/corr/abs-2402-01912 ; lacombe-etal-2024-parler-tts proposes large-scale annotation of basic style tags (e.g., pitch, speed) by leveraging signal processing techniques and rule-based binning. Other efforts, such as FleSpeech DBLP:journals/corr/abs-2501-04644 and VccmDataset DBLP:journals/corr/abs-2406-01205 , re-caption existing speech datasets like TextrolSpeech using large language models to improve label quality and expressiveness. VoxInstruct DBLP:conf/mm/0002QJZLZ0024 and SpeechCraft DBLP:conf/mm/Jin0W0ZZQ024 further extend this direction by introducing more fine-grained emotion tags. In addition, AudioBox DBLP:journals/corr/abs-2312-15821 integrates multiple strategies, combining large-scale automatically generated basic tags with high-quality human-annotated rich style labels to build a scalable and diverse dataset. It also supports sound events such as laughter and clapping, which are important for expressive speech generation.
@@
Table 1 summarizes existing English style-captioned datasets. Please refer to Appendix A.1 for the description of each speech style trait. To the best of our knowledge, the proposed CapSpeech is the first large-scale open-source dataset that covers a wide range of both intrinsic speaker traits and expressive speaking styles. Similar to AudioBox DBLP:journals/corr/abs-2312-15821 , CapSpeech also supports speech generation with sound events.
2.2 Speech Style-captioned TTS Models
@@
Most existing methods for speech style-captioned TTS build upon prior audio-prompted TTS frameworks. For example, PromptTTS DBLP:conf/icassp/GuoLWZT23 is developed on top of FastSpeech 2 DBLP:conf/iclr/0006H0QZZL21 , while PromptTTS2 DBLP:conf/iclr/LengGSJ0LLYZS0024 is based on NaturalSpeech 2 DBLP:conf/iclr/ShenJ0LL00Z024 . Similarly, Salle DBLP:conf/icassp/JiZ00CDHZ24 builds on VALL-E DBLP:journals/corr/abs-2301-02111 , and Parler-TTS DBLP:journals/corr/abs-2402-01912 further introduces cross-attention mechanisms to better integrate caption features. In parallel, diffusion-based approaches such as InstructTTS DBLP:journals/taslp/YangLHWM24 and AudioBox DBLP:journals/corr/abs-2312-15821 have also emerged, offering alternative generative paradigms for style-controllable speech synthesis. In this work, we adopt both autoregressive and non-autoregressive models to evaluate performance on our proposed benchmark.
3 CapSpeech Benchmark
@@
3.1 Tasks and Applications
As shown in Figure 1, the CapSpeech benchmark includes the following five tasks:
@@ 3.1 Tasks and Applications
EmoCapTTS generates speech from text input conditioned on a natural language caption that describes both the speaker’s emotional state and identity. Unlike traditional emotion TTS systems that rely on discrete categories (e.g., happy, sad, angry) DBLP:conf/icassp/Tang [^0] W24 ; DBLP:conf/interspeech/Tang0W0023 , EmoCapTTS enables more flexible and expressive emotion control through free-form textual descriptions. This subtask balances multi-speaker control with basic emotion expression, targeting applications like story narration and gaming NPCs where multiple AI speakers are required.
@@
      Pretraining       Emilia-en, GigaSpeech, MLS-en, CommonVoice,
@@
      CapTTS, EmoCapTTS, AccCapTTS       LibriTTS-R, EARS, Expresso,
      VCTK, VoxCeleb, VoxCeleb2
@@
3.2 Datasets
In this section, we introduce the data collection process in CapSpeech [^1] https://huggingface.co/datasets/OpenSound/CapSpeech and summarize the data sources used across different tasks in Table 2.
3.2.1 Pretraining Sets
@@
For CapTTS-PT, we generate captions on four English speech sources: the English portion of Emilia DBLP:conf/slt/HeSWLGHLYLSWCZW24 , the English portion of Multilingual LibriSpeech (MLS) DBLP:conf/interspeech/PratapXSSC20 , GigaSpeech DBLP:conf/interspeech/ChenCWDZWSPTZJK21 , and CommonVoice DBLP:conf/lrec/ArdilaBDKMHMSTW20 . We apply data cleaning and annotate age, gender, pitch, expressiveness of tone, and speaking rate for MLS, GigaSpeech, and CommonVoice, and use an LLM [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 to generate natural language captions based on these traits. For Emilia, we adopt the style annotations provided in ParaSpeechCaps DBLP:journals/corr/abs-2503-04713 , which include 59 diverse style tags. Please refer to Appendix A.2 for details.
@@
For CapTTS-SE-PT, we simulated data using the LibriTTS-R speech corpus and three sound event corpora: VGGSound DBLP:conf/icassp/ChenXVZ20 , FSDKaggle2018 DBLP:conf/dcase/FonsecaPFEFPS18 , and ESC-50 piczak2015dataset . We refer to this dataset as CapSpeech-PT-SEDB. We curate a collection of 394 distinct sound events drawn from the categories of these datasets. sound events are introduced into speech using two modes: (i) insertion, where the event is inserted at a specific location; and (ii) background, where the event is layered beneath the speech. For the background mode, we use noisy audio from VGGSound, while the insertion mode relies on cleaner sources from FSDKaggle2018 and ESC-50. For each LibriTTS-R sample, a sound event and mode are randomly selected. Forced alignment is performed using the Montreal Forced Aligner [^3] https://huggingface.co/datasets/cdminix/libritts-r-aligned to determine insertion points—either at the beginning, middle, or end of the utterance. The speech and sound events are then mixed at a random SNR ranging from –3 dB to 6 dB. To scale up the data, each speech sample is simulated five times with different configurations. See Appendix A.3 for details. In total, the dataset contains 10,054,530 speech–caption pairs, with 7,894 pairs reserved for validation and 7,959 pairs held out for testing.
3.2.2 SFT Sets for CapTTS, EmoCapTTS, and AccCapTTS

### v2

@@ I Introduction
In recent years, large-scale text-to-speech (TTS) synthesis has witnessed remarkable progress, exemplified by models such as VALL-E 2 [1], NaturalSpeech 3 [2], LLasa [3], SimpleSpeech 2 [4], CosyVoice 2 [5], and MaskGCT [6]. Most of these efforts have centered on modeling audio-based speaker characteristics, such as zero-shot TTS with audio prompts [7, 8, 9], category-based traits [10, 11, 12, 13], or embedding-based speaker representations [14, 15]. However, the detailed understanding of speaking style from audio, particularly its subtle nuances, has received limited attention.
@@
The style of speech encompasses both intrinsic traits tied to a speaker’s identity (e.g., age, gender, timbre) and expressive style traits specific to individual utterances (e.g., emotion, speaking rate). Recent studies have introduced the use of natural language captions to describe these stylistic elements—a paradigm referred to as prompt TTS, expressive TTS, or style-captioned TTS [16, 17, 18, 19, 20, 21, 22]. In this work, we adopt the term style-captioned TTS and abbreviate it as CapTTS.
@@
Developing a CapTTS system necessitates a large corpus of audio-caption pairs accompanied by transcriptions, the annotation of which is labor-intensive and costly. Parler-TTS [23, 24] addressed this challenge by automatically annotating basic speech style attributes such as pitch and speed using signal processing tools. ParaSpeechCaps [25] collected speaker-level intrinsic tags and utterance-level situational tags, LibriTTS-P [26] offered speaker identity-based captions, and EmoVoice-DB [27] provided emotion-based captions. However, these existing datasets lack a unified and comprehensive framework for style captioning, making cross-domain comparisons difficult. Moreover, there has been limited exploration of downstream applications, such as transferring models to new caption styles or incorporating sound events into the synthesized speech.
@@
To address the challenges outlined above, we present CapSpeech, a novel benchmark featuring standardized and comprehensive datasets for CapTTS and its related downstream tasks. CapSpeech comprises a pretraining stage with large-scale captioned speech, as well as five downstream tasks: CapTTS, CapTTS-SE, accent-captioned TTS (AccCapTTS), emotion-captioned TTS (EmoCapTTS), and TTS for chat agent (AgentTTS). The pretraining datasets consist of over 10 million machine-annotated audio-caption pairs, while the downstream datasets contain 358,783 human-annotated audio-caption pairs. These datasets encompass a broad range of intrinsic speaker traits and expressive style traits, curated from a wide array of audio sources, including Emilia [28], GigaSpeech [29], CommonVoice [30], MLS [31], LibriTTS-R [32], VoxCeleb [33], VoxCeleb2 [34], EARS [35], Expresso [20], VCTK [36], VGGSound [37], FSDKaggle2018 [38], and ESC-50 [39]. In addition, we introduce two new datasets: one for AgentTTS, built using professionally recorded voice actor speech, and another for CapTTS-SE, processed by five experienced audio engineers.
@@
In addition, we develop two style-captioned TTS models based on state-of-the-art generative TTS backbones: one autoregressive (AR) and one non-autoregressive (NAR). We evaluate these models on the CapSpeech benchmark across 5 downstream tasks. For a comprehensive evaluation, both objective and subjective metrics are employed to assess speech style consistency, audio quality, and intelligibility.
@@
  • We introduce CapSpeech, a large-scale style-captioned benchmark for CapTTS that encompasses diverse speech styles.
  • We propose 5 CapTTS-related downstream tasks, 4 of which are new, aimed at practical CapTTS applications, introduce 9,353,823 new captions for existing speech corpora, and curate 3 new speech datasets specifically designed for the CapTTS-SE and AgentTTS tasks, thereby enhancing the benchmark’s coverage of real-world scenarios.
  • We train and evaluate two CapTTS models, CapSpeech-AR and CapSpeech-NAR, which address the lack of benchmarks for CapTTS downstream tasks, demonstrate the effectiveness of the proposed dataset, and highlight key challenges and future research directions.
  • We publicly release our datasets, listening samples, source code, pretrained checkpoints, and evaluation toolkit to support future research [^1] https://github.com/WangHelin1997/CapSpeech. All resources are released under the CC BY-NC 4.0 license (Creative Commons Attribution-NonCommercial), which permits use for non-commercial research purposes with appropriate attribution.
TABLE I: A comparison of English speech style-captioned datasets. I1–I5 denote intrinsic speaker traits: age (I1), gender (I2), timbre (I3), mean pitch (I4), and accent (I5). E1–E4 represent expressive style traits: speaking rate (E1), emotion (E2), expressiveness of tone (E3), and volume (E4).
@@
PromptSpeech [16] 0.3k
Expresso [20] 47
EARS [35] 60
TextrolSpeech [40] 0.3k
SpeechCraft [17] 2.4k
VccmDataset [41] 0.3k
LibriTTS-P [26] 0.6k
DreamVoiceDB [42] 0.3k
ParlerTTS [23] 45k
ParaSpeechCaps [25] 2.9k
NonVerbalSpeech-38K [43] 0.1k
PromptTTS2 [18] 44k
VoxInstruct [44] 1.5k
FleSpeech [45] 0.6k
Audiobox [46] >0.5k
@@
II Related Works
II-A Speech Style-Captioned Datasets
@@
Speech datasets with natural language style captions offer greater flexibility and controllability compared to those that rely on audio prompts. Early efforts such as NLSpeech [21] recruited human annotators to describe speech emotions, laying the groundwork for later datasets like PromptStyle [47] and MM-TTS [48]. TextrolSpeech [40] aggregates several existing emotion datasets, primarily focusing on emotional states and a few basic style tags. In contrast, Expresso [20] and EARS [35] broaden the scope by covering a wider range of spontaneous expressive styles. Meanwhile, datasets such as LibriTTS-P [26] and DreamVoiceDB [42] focus on intrinsic speaker traits, providing captioned annotations for LibriTTS-R [32], while Coco-Nut [49, 50] captures subjective descriptions of voice characteristics beyond traditional acoustic parameters. More recently, ParaSpeechCaps [25] introduces a human-annotated dataset that covers a richer and more diverse set of speaking style attributes. Given the high cost and limited scalability of human annotation, several works have explored automatic or semi-automatic approaches to expand style-captioned speech datasets. PromptSpeech [16] and PromptTTS2 [18] synthesize speech with diverse speaker identities and emotions using commercial TTS systems. Parler-TTS [23, 24] proposes large-scale annotation of basic style tags (e.g., pitch, speed) by leveraging signal processing techniques and rule-based binning. Other efforts, such as FleSpeech [45] and VccmDataset [41], re-caption existing speech datasets like TextrolSpeech using large language models to improve label quality and expressiveness. VoxInstruct [44] and SpeechCraft [17] further extend this direction by introducing more fine-grained emotion tags. In addition, AudioBox [46] integrates multiple strategies, combining large-scale automatically generated basic tags with high-quality human-annotated rich style labels to build a scalable and diverse dataset. It also supports sound events such as laughter and clapping, which are important for expressive speech generation.
@@
Table I summarizes existing English style-captioned datasets. To the best of our knowledge, the proposed CapSpeech is the first large-scale open-source dataset that covers a wide range of both intrinsic speaker traits and expressive speaking styles. Similar to AudioBox [46] and NonVerbalSpeech-38K [43], CapSpeech also supports speech generation with sound events.
II-B Speech Style-captioned TTS Models
@@
Most existing methods for speech style-captioned TTS build upon prior audio-prompted TTS frameworks. For example, PromptTTS [16] is developed on top of FastSpeech 2 [51], while PromptTTS2 [18] is based on NaturalSpeech 2 [52]. Similarly, Salle [40] builds on VALL-E [53], and Parler-TTS [23] further introduces cross-attention mechanisms to better integrate caption features. In parallel, diffusion-based approaches such as InstructTTS [21] and AudioBox [46] have also emerged, offering alternative generative paradigms for style-controllable speech synthesis. In this work, we adopt both autoregressive and non-autoregressive models to evaluate performance on our proposed benchmark.
III CapSpeech Benchmark
@@
III-A Tasks and Applications
As shown in Fig. 1, the CapSpeech benchmark includes the following five tasks:
@@ III-A Tasks and Applications
EmoCapTTS generates speech from text input conditioned on a natural language caption that describes both the speaker’s emotional state and identity. Unlike traditional emotion TTS systems that rely on discrete categories (e.g., happy, sad, angry) [15, 54], EmoCapTTS enables more flexible and expressive emotion control through free-form textual descriptions. This subtask balances multi-speaker control with basic emotion expression, targeting applications like story narration and gaming NPCs where multiple AI speakers are required.
@@
III-B Datasets
In this section, we introduce the data collection process in CapSpeech [^2] https://huggingface.co/datasets/OpenSound/CapSpeech and summarize the data sources used across different tasks in Table II. The total amount of annotated data is 30.8k hours for age, 33.5k hours for gender, 0.4k hours for timbre, 33.5 k hours for pitch, 2.5 k hours for accent, 33.5 k hours for speaking rate, 2.4 k hours for emotion, 33.5 k hours for expressiveness of tone, and 2.7 k hours for volume.
@@ III-B Datasets
Pretraining Emilia-en, GigaSpeech, MLS-en,
@@
CapTTS, EmoCapTTS, LibriTTS-R, EARS, Expresso,
AccCapTTS VCTK, VoxCeleb, VoxCeleb2
@@
III-B1 Pretraining Sets
@@
CapTTS-PT comprises four English speech corpora: the English portion of Emilia [28], the English portion of Multilingual LibriSpeech (MLS) [31], GigaSpeech [29], and CommonVoice [30]. Since the overall audio quality of CommonVoice is poor, we only include the age groups under 20 and over 65, which are not well covered in other datasets. These datasets cover a wide range of speech sources: LibriVox, YouTube, Podcasts, Audiobooks, and etc. For Emilia, we directly use the style annotations provided in ParaSpeechCaps [25], which include 59 diverse style tags. In total, there are 9,053,734 speech–caption pairs.
We annotate age, gender, pitch, expressiveness of tone, and speaking rate for MLS, GigaSpeech, and CommonVoice, and use an LLM [^3] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 to generate natural language captions based on these traits. Althouth these datasets provide transcriptions, there are some audio files with noisy labels or multiple speakers. Thus, we apply three data cleaning steps: (i) samples with durations between 3 and 18 seconds are selected. (ii) we filter out samples with word error rates (WER) greater than $25\%$, estimated using the OpenAI Whisper toolkit [^4] https://huggingface.co/openai/whisper-large-v3-turbo [55]; (iii) we remove noisy utterances with estimated signal-to-noise ratios (SNR) below 20 dB, computed using the Squim Objective metric [56]. We use a pre-trained age and gender estimator [^5] https://github.com/audeering/w2v2-age-gender-how-to [57] to annotate speaker demographics, if they are not provided by the dataset creator. The age groups are divided into five categories: "child" (1–12 years), "teenager" (13–19 years), "young adult" (20–39 years), "middle-aged adult" (40–64 years), and "elderly" (65 years and older). Following Parler-TTS [23], pitch and expressiveness are measured using speaker-level mean and utterance-level standard deviation of pitch, computed with the PENN library [^6] https://github.com/interactiveaudiolab/penn. The speaker-level mean is used to generate a label for speaker pitch relative to gender, and the standard deviation is used as a proxy for how monotone or animated an individual utterance is. Speaking rate is calculated by dividing the number of phonemes in the transcription by the total duration of the silence-removed utterance.
Next, we convert all the variables described above into natural language sentences. To do so, we first generate keyword representations for each variable. Discrete labels provided by the dataset creators, such as gender, are used directly as keywords. For continuous variables like pitch and speaking rate, we apply binning to map them into discrete categories. After binning all continuous variables, we obtain a complete set of keywords for gender, accent, pitch, expressiveness of tone, and speaking rate.
More specifically, the age category includes "child" ($0.15\%$), "teenager" ($0.20\%$), "young adult" ($32.72\%$), "middle-aged adult" ($63.44\%$), and "elderly" ($3.49\%$); the gender category includes "male" ($57.48\%$) and "female" ($42.52\%$); the pitch category includes "very low-pitch" ($3.22\%$), "low-pitch" ($7.19\%$), "slightly low-pitch" ($26.26\%$), "moderate pitch" ($28.05\%$), "slightly high-pitch" ($21.34\%$), "high-pitch" ($12.45\%$) and "very high-pitch" ($1.49\%$); the expressiveness of tone category includes "very monotone" ($10.59\%$), "monotone" ($48.56\%$), "slightly expressive and animated" ($24.58\%$), "expressive and animated" ($8.32\%$) and "very expressive and animated" ($7.96\%$); the speaking rate category includes "very slowly" ($0.01\%$), "slowly" ($1.60\%$), "slightly slowly" ($15.03\%$), "moderate speed" ($59.81\%$), "slightly fast"($18.98\%$), "fast" ($4.10\%$) and "very fast" ($0.45\%$).
We then apply the LLM to generate natural language captions based on these keywords. To guide the model, we construct a prompt and curate a diverse set of manually verified examples. During caption generation, random examples are paired with different audio files to enhance the diversity of expressions. Finally, we apply post-processing to remove excessively long captions. Below is the LLM prompts used:
Prompt Design for Pretraining Sets System: You are a voice style caption generator. User: Please generate a single text description of a speech sample using the provided keywords. Definition of Keywords:
• Gender (e.g., male, female)
Keywords: Gender: [e.g., female]; Age: [e.g., teenager]; Tone: [e.g., bright, expressive]; Pace: [e.g., quick]; Pitch: [e.g., high].
LLM Assistant: [e.g., A teenage girl speaks with a high-pitched, bright voice that’s lively and expressive, delivering her words at a quick pace.]
To assess potential biases introduced by machine annotation, we conducted a systematic human evaluation on the CapTTS-PT dataset. We randomly sampled 500 examples, each annotated independently by two experienced audio engineers. For each example, individual tags (e.g., age, gender, speaking rate) were evaluated with a binary correctness label (0 or 1) indicating whether the machine-predicted tag accurately reflected the speech. Also, a caption-level quality score was assigned on a 1–5 Likert scale to assess the overall coherence, coverage, and naturalness of the generated captions. Table III shows the detailed results, reporting the mean values along with their $95\%$ confidence intervals. We can see that the average per-tag correctness is $95.5\%$, demonstrating reliable labeling. The average caption quality score is 4.2 out of 5, indicating strong overall fidelity.
TABLE III: Human evaluation on the CapTTS-PT dataset.
Attribute Score
Age 0.979 ± 0.019
Gender 1.000 ± 0.000
Pitch 0.944 ± 0.030
Speaking rate 0.991 ± 0.012
Expressive of tone 0.863 ± 0.004
Caption 4.220 ± 0.071
@@
For CapTTS-SE-PT, we simulated data using the LibriTTS-R speech corpus and three sound event corpora: VGGSound [37], FSDKaggle2018 [38], and ESC-50 [39]. We curate a collection of 394 distinct sound events drawn from the categories of these datasets. We refer to this dataset as CapSpeech-PT-SEDB. Sound events are introduced into speech using two modes: (i) insertion, where the event is inserted at a specific location; and (ii) background, where the event is layered beneath the speech. For the background mode, we use noisy audio from VGGSound, while the insertion mode relies on cleaner sources from FSDKaggle2018 and ESC-50. For each LibriTTS-R sample, a sound event and mode are randomly selected. Forced alignment is performed using the Montreal Forced Aligner [^7] https://huggingface.co/datasets/cdminix/libritts-r-aligned to determine insertion points—either at the beginning, middle, or end of the utterance. To ensure the simulated audio sounds natural and smooth, we select insertion points between words with a minimum interval of 0.3 seconds, ensuring a pause that does not disrupt word continuity. The speech and sound event are then mixed at a random SNR ranging from –3 dB to 6 dB. To scale up the data, each speech sample is simulated five times with different configurations.
TABLE IV: Sound event examples in CapSpeech.
Applause Bark Burping or eructation Bus
Chime Writing Drawer open or close Cough
Fart Finger snapping Fireworks Crow
Keys jangling Knock Laughter Meow
Microwave oven Clock tick Shatter Hand saw
Squeak Tearing Telephone Computer
keyboard Dog Rooster Rain
Crying baby Door knock Scissors Helicopter
Sea waves Sneezing Mouse click Chainsaw
Pig Crackling fire Clapping Keyboard typing
Siren Cow Crickets Breathing
Door wood creaks Car horn Frog Chirping birds
Coughing Can opening Engine Cat
Water drops Footsteps Washing machine Hand saw
Train Hen Wind Laughing
Vacuum cleaner Church bells Insects (flying) Pouring water
Brushing teeth Clock alarm Airplane Sheep
Toilet flush Snoring Gunshot or gunfire Thunderstorm
Glass breaking Drinking, sipping
In total, the dataset contains 10,054,530 speech–caption pairs, with 7,894 pairs reserved for validation and 7,959 pairs held out for testing.
III-B2 SFT Sets for CapTTS, EmoCapTTS, and AccCapTTS